# Limit theorems for non-Hermitian ensembles

### v2

@@ Abstract.
The distribution of the modulus of the extreme eigenvalues is investigated for the complex Ginibre and complex induced Ginibre ensembles in the limit of large dimensions of random matrices. The limiting distribution of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble, with a proportional rectangularity index, is the Gumbel distribution. The independence of these extrema is established, at appropriate scaling, for large matrices from the complex Ginibre ensemble as well as from the complex induced Ginibre ensemble for either fixed or proportional rectangularity indexes. The joint distribution of the eigenvalue moduli for the complex Ginibre ensemble is derived in this paper. It is equal to the joint distribution of independent random variables, each following a Gamma-Rayleigh distribution. In the limit of a large size of the complex Ginibre matrices, the left and right tail distributions of the minimum modulus are the Rayleigh and Weibull distributions, respectively. The limiting left tail distribution of minimum modulus is the same for these non-Hermitian ensembles when the rectangularity index of the complex induced Ginibre ensemble is equal to zero. This phenomenon is also verified for the right tail distribution of this minimum.
@@ 1. Introduction
$P(J)=\frac{1}{\pi^{N^{2}}}\exp\left(-\operatorname{Tr}\left(JJ^{*}\right)\right)$
@@
$P_{N}(z)=\frac{1}{\pi^{N}C_{N}}\exp\left(-\sum_{k=1}^{N}|z_{k}|^{2}\right)|\Delta(z)|^{2}$
@@
$C_{N}=\frac{1}{\pi^{N}}\int_{\mathbb{C}^{N}}\exp\left(-\sum_{k=1}^{N}|z_{k}|^{2}\right)\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{2}\prod_{k=1}^{N}d^{2}z_{k}=\prod_{k=1}^{N}\Gamma(k+1)$
@@
Section 2 is devoted to the analysis of the distribution of the eigenvalue moduli for the complex Ginibre ensemble. The main results are derived for the spectral radius and the minimum modulus. The limiting left and right tail distributions of the minimum modulus is also studied for large size of matrices in Section 2.1. Pursuing the analysis in the scaling limit $\sqrt{N}$ for this random matrix ensemble as $N$ goes to infinity, the independence of the scaled minimum moduli with respect to the scaled spectral radius is established in Section 2.3. The eigenvalues of matrices from the complex Ginibre ensemble form a determinantal point process and the joint probability density of their respective radius has been investigated in the work of E. Kostlan [18] extended in [16]. This result is also discussed in the present paper in Section 2.2.
@@
$P_{FZ}(A)\propto\exp\left(-\operatorname{Tr}V(A^{\dagger}A)\right)$
@@
$P_{indG}^{(\beta)}(G)=\Gamma^{(\beta)}\left(\det G^{\dagger}G\right)^{\frac{\beta}{2}L}\exp\left(-\operatorname{Tr}G^{\dagger}G\right)\\$
@@
$\Gamma^{(\beta)}=\pi^{-\frac{\beta}{2}N^{2}}\left(\frac{\beta}{2}\right)^{\frac{1}{2}N(N-L)}\prod_{j=1}^{N}\frac{\Gamma(\frac{\beta}{2}j)}{\Gamma(\frac{\beta}{2}(j+L)}$
@@
$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{\Gamma_{N}}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\\$
@@
$\Gamma_{N}$ $=\int\cdots\int\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\prod_{j=1}^{N}d^{2}\lambda_{j}$
$=N!\pi^{N}\prod_{k=1}^{N}\Gamma(k+L)$
@@ 2. Limit theorems for the complex Ginibre ensemble
where each random variables $\gamma_{k}$ are independent and follows a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$ , with $\alpha_{k}=k$ and $\delta_{k}=1$. The Gamma-Rayleigh distribution is derived in reference [1].
Corollary 3.
Let $\left((r_{i}^{(N)}(A))^{2}\right)_{i=1}^{N}$ denote the squared (unordered) moduli of the $N$ eigenvalues $\left(z_{i}^{(N)}(A)\right)_{i=1}^{N}$ of a $N\times N$ complex Ginibre matrix $A$. The distribution of the squared spectral radius $\left(r_{max}^{(N)}(A)\right)^{2}$ is the joint distribution of the elements of a finite sequence $\{\zeta_{k}\}_{k=1}^{N}$. The $\zeta_{k}$ are independent random variables. Each $\zeta_{k}$ follows a Gamma distribution with shape parameter $\alpha_{k}=k$ and scale parameter $\delta_{k}=1$, i.e., $\zeta_{k}\sim Gamma(k,1)$.
Proof.
A simple proof of this result is as follows. Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble. Let $\left(r_{i}^{(N)}(A)\right)_{1\leq i\leq N}$ denote the moduli of the eigenvalues of the matrix $A$. The joint distribution of the square moduli $\left((r_{i}^{(N)}(A))^{2}\right)_{i=1}^{N}$ is derived from Equation (4) and is expressed as
$P\left((r_{max}^{(N)}(A))^{2}\leq r\right)=\prod_{k=1}^{N}\frac{\gamma(k,r)}{\Gamma(k)}=\prod_{k=1}^{N}P(\zeta_{k}\leq r)\\$
where $P(\zeta_{k}\leq r)=\frac{\gamma(k,r)}{\Gamma(k)}$ is the cumulative distribution function of the random variable $\zeta_{k}$ following a Gamma distribution with shape parameter $\alpha_{k}=k$ and rate parameter $\delta_{k}=1$.
The corresponding probability density function of the largest modulus $r_{max}^{(N)}(A)$ is derived from Equation (4). Taking the logarithm of the cumulative distribution function $P(r_{max}^{(N)}(A)\leq r)$ and then the first derivative with respect to $r$ of $\log\left(P(r_{max}^{(N)}(A)\leq r)\right)$, the probability density funvtion of the spectral radius $r_{max}^{(N)}(A)$ is,
@@ 2.1. Limiting left and right tail distributions of the minimum modulus as $N$ goes to infinity
Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble. The left tail distribution of minimum modulus $r_{min}^{(N)}(A)$ (i.e., the smallest values of $r_{min}^{(N)}(A)$) converges to the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$ as $N$ goes to infinity.
More precisely,
$\lim_{r\longrightarrow 0}\lim_{N\longrightarrow+\infty}P(r_{min}^{(N)}(A)<r)=1-e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (9)
@@
2.2. The joint distribution of the eigenvalue moduli for the complex Ginibre ensemble
G. Dubach proved in [6] that the collection of squared moduli of the eigenvalues of $N\times N$ matrices from the complex Ginibre ensemble has the same distribution as the collection of independent random variables $\zeta_{k}$, each following a Gamma distribution with shape parameter $\alpha_{k}=k$ and scale parameter $\delta_{k}=1$, i.e., $\zeta_{k}\sim Gamma(k,1)$. He did provide proof of this result which involves the use of the Andreief’s integration formula [9] with radially symmetric functions defined on the complex plane.
Theorem 6.
The joint distribution of the unordered eigenvalue moduli $\{r_{k}^{(N)}(A)\}_{k=1}^{N}$ of $N\times N$ matrices $A$ from the complex Ginibre ensemble is the joint distribution of independent random variables $\zeta_{k}$ each following a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$, with $\alpha_{k}=k$ and $\delta_{k}=1$,
$\{r_{1}^{(N)}(A),...,r_{N}^{(N)}(A)\}\overset{d}{=}\{\zeta_{1},...,\zeta_{N}\}$
Proof.
Let $\left(\Omega,\mathcal{F},\mu\right)$ denote a measure space with $\Omega=\mathbb{C}$, $\mathcal{F}$ a $\sigma$-algebra on the complex plane and the measure $\mu$, a nonnegative function, defined on the measurable space $\left(\Omega,\mathcal{F}\right)$. Let $Q$ denote the set of $N$ moduli $r_{k}^{(N)}(A)$ of $N\times N$ matrices $A$ from the complex Ginibre ensemble, i.e., $Q=\left(q_{k}\right)_{k=1}^{N}=\left(r_{k}^{(N)}(A)\right)_{k=1}^{N}$. Let $\varphi$ denote a (suitable) test function with $\mu$ the reference measure on the complex plane. Using the Andreief’s integration formula
$\mathbb{E}\left[\prod_{q_{k}\in Q}\varphi(q_{k})\right]$ $=\int...\int\prod_{k=1}^{N}\varphi(|z_{k}|)P_{N}(z_{1},...,z_{N})\prod_{k=1}^{N}d\mu(z_{k})$
$=\prod_{k=1}^{N}\int_{0}^{+\infty}\varphi(r)\frac{2r^{2k-1}e^{-r^{2}}}{\Gamma(k)}dr=\prod_{k=1}^{N}\int_{0}^{+\infty}\varphi(r)p_{GR(k,1)}(r)dr=\prod_{k=1}^{N}\mathbb{E}\left[\varphi(\zeta_{k})\right]$
where each $\zeta_{k}$ follows a Gamma-Rayleigh distribution with shape parameter $\alpha_{k}=k$ and scale parameter $\delta_{k}=1$. The Gamma-Rayleigh distribution is derived in the work of E. Akarawak, I. Adeleke and R. Okafor [1].
Remark 2.2.
[…]
$\left\{|z_{1}|,...,|z_{n}|\right\}\overset{d}{=}\left\{R_{(1)},...,R_{(n)}\right\}$
where $|z_{1}|\geq...\geq|z_{n}|$ are the order statistics of the moduli and where $R_{(1)}\geq...\geq R_{(n)}$ is the order statistics of independent random variables $R_{1},...,R_{n}$ with $R_{k}$ of density proportional to $t\longmapsto t^{2k-1}e^{-nV(t)}1_{[0,+\infty)}(t)$ with $V(t)=t^{2}$, for every $1\leq k\leq n$.
Remark 2.3.
The joint distribution of the moduli of $N\times N$ matrices for the complex Ginibre ensemble has been investigated in the work of E. Kostlan [18]. He stated that the collection of moduli of the eigenvalues of $N\times N$ matrices from the complex Ginibre ensemble has the same distribution as the collection of random variables $\{X_{2k}\}_{1\leq k\leq N}$ where, $\forall k\in\{1,...,N\}$, $X_{2k}$ follows a chi distribution $\chi_{2k}$. A slight extension of this result would correspond to state the next Theorem 7.
Theorem 7.
The collection of unordered eigenvalue moduli $|z_{k}|$ of $N\times N$ matrices from the complex Ginibre ensemble has the same distribution as the collection of random variables $\{\frac{X_{2k}}{\sqrt{2}}\}_{1\leq k\leq N}$ where, $\forall k\in\{1,...,N\}$, $X_{2k}$ follows a chi distribution $\chi_{2k}$.
Proof.
Let $Q$ denote the set of $N$ moduli $|z_{k}|$ of matrices from the complex Ginibre ensemble, i.e., $Q=\left(q_{k})\right)_{k=1}^{N}$. Let $f$ denote a (suitable) test function. Using the Andreief’s integration formula,
$\mathbb{E}\left[\prod_{q_{k}\in Q}f(q_{k})\right]$ $=\int...\int\prod_{k=1}^{N}f(|z_{k}|)P_{N}(z_{1},...,z_{N})\prod_{k=1}^{N}d\mu(z_{k})$
$=\prod_{k=1}^{N}\int_{0}^{+\infty}f(v/\sqrt{2})\frac{v^{2k-1}e^{-v^{2}/2}}{2^{k-1}\Gamma(k)}dv=\prod_{k=1}^{N}\int_{0}^{+\infty}f(v/\sqrt{2})p_{\chi_{2k}}(v)dv=\prod_{k=1}^{N}\mathbb{E}\left[f\left(\frac{\alpha_{k}}{\sqrt{2}}\right)\right]$
Each random variable $\alpha_{k}$ follows a chi distribution with parameter $2k$, i.e., $\alpha_{k}\sim\chi_{2k}$ for $k\in\{1,...,N\}$.
@@ 2.2. The joint distribution of the eigenvalue moduli for the complex Ginibre ensemble
$\lim_{N\rightarrow+\infty}P\left(R_{N}\leq R^{(N,x)}\right)=\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\gamma(k+1,N\left(R^{(N,x)}\right)^{2})}{\Gamma(k+1)}=F_{X}(x)=e^{-e^{-x}}$
@@
$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=E_{0}^{(2)}\left(\sqrt{N}r\right)$
@@ 3.1. Limit theorems of the outer and inner edges of the ring
Let $A$ denote a $N\times N$ random matrix from the complex Ginibre ensemble. The limiting left tail probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is the probability density function of the Rayleigh distribution with parameter $\frac{1}{\sqrt{2}}$, in the neighbourhood of zero, as $N$ goes to infinity. More precisely,
$\lim_{r\longrightarrow 0}\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (25)
@@
The limit of partial product $\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$, as $N$ goes to infinity, has been investigated in Section 2 for the complex Ginibre ensemble, and for small $r$ in the neighbourhood of zero, it is expressed as
@@
The limit of the partial sum is derived as follows. Knowing that, for $r$ in the neighbourhood of zero, $\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$,
@@
Thus, the limiting left tail probability density function of the minimum of moduli for matrices from the complex Ginibre ensemble is
@@
The minimum modulus of a $N\times N$ matrix $G$ from the complex induced Ginibre ensemble, with rectangularity index $L=1$, has a left tail probability density function of the Generalised Gamma distribution with scale parameter $\alpha=1$ and other parameters $\kappa=2$ and $\gamma=4$. The corresponding limiting probability density function, for $r$ in the neighbourhood of zero, as $N$ goes to infinity, is
$\lim_{r\longrightarrow 0}\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=2r^{3}e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (26)
@@
Also, the upper incomplete Gamma function $\Gamma(k,x)$, for $x$ approaches zero and $k\in\mathbb{N}\setminus\{0\}$, is $\Gamma(k,x)=\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$.
For $r$ close to zero, $\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$. This implies that the limit of the partial sum, as $r$ approaches zero and as $N$ goes to infinity, is
@@
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L\geq 0$. The left tail distribution of minimum of the moduli $r_{min}^{(N)}(G)$ is the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$ in the limit as $N$ goes to infinity and $r$ approaches zero zero,
@@
This result is derived using the same methodological approach presented in section 2.1. For positive values of $r$ in the neighbourhood of zero,
@@
Setting the rectangularity index $L$ to zero, the limiting distribution of the minimum modulus $r_{min}^{(N)}(G)$ for the complex induced Ginibre ensemble corresponds to the limiting distribution of the minimum modulus $r_{min}^{(N)}(A)$ for the complex Ginibre ensemble, i.e., for small $r$
@@
For positive values of $r$ in the neighbourhood of zero, as $N$ goes to infinity, and for a rectangularity index $L=0$,
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)+O\left(r^{5}\right)$

### v3

@@ Abstract.
The distribution of the modulus of the extreme eigenvalues is investigated for the complex Ginibre and complex induced Ginibre ensembles in the limit of large dimensions of random matrices. The limiting distribution of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble, with a proportional rectangularity index, is the Gumbel distribution. The independence of these extrema is established, at appropriate scaling, for large matrices from the complex Ginibre ensemble as well as from the complex induced Ginibre ensemble for fixed and proportional rectangularity indexes. In the limit of a large size of the complex Ginibre matrices, the left and right tail distributions of the minimum modulus are the Rayleigh and Weibull distributions, respectively. The limiting left tail distribution of the minimum modulus is the same for these non-Hermitian ensembles when the rectangularity index of the complex induced Ginibre ensemble is equal to zero. This phenomenon is also verified for the right tail distribution of this minimum.
@@ 1. Introduction
$P(J)=\frac{1}{\pi^{N^{2}}}\exp\left(-\operatorname{Tr}\left(JJ^{*}\right)\right)$ (2)
@@
$P_{N}(z)=\frac{1}{\pi^{N}C_{N}}\exp\left(-\sum_{k=1}^{N}|z_{k}|^{2}\right)|\Delta(z)|^{2}$ (3)
@@
$C_{N}=\frac{1}{\pi^{N}}\int_{\mathbb{C}^{N}}\exp\left(-\sum_{k=1}^{N}|z_{k}|^{2}\right)\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{2}\prod_{k=1}^{N}d^{2}z_{k}=\prod_{k=1}^{N}\Gamma(k+1)$ (4)
@@
Section 2 is devoted to the analysis of the distribution of the eigenvalue moduli for the complex Ginibre ensemble. The main results are derived for the spectral radius and the minimum modulus. The limiting left and right tail distributions of the minimum modulus is also studied for large size of matrices in Section 2.1. Pursuing the analysis in the scaling limit $\sqrt{N}$ for this random matrix ensemble as $N$ goes to infinity, the independence of the scaled minimum moduli with respect to the scaled spectral radius is established in Section 2.2. The eigenvalues of matrices from the complex Ginibre ensemble form a determinantal point process and the joint probability density of their respective radius has been investigated in [16], and is extended in [14].
@@
$P_{FZ}(A)\propto\exp\left(-\operatorname{Tr}V(A^{\dagger}A)\right)$ (5)
@@
$P_{indG}^{(\beta)}(G)=\Gamma^{(\beta)}\left(\det G^{\dagger}G\right)^{\frac{\beta}{2}L}\exp\left(-\operatorname{Tr}G^{\dagger}G\right)\\$ (6)
@@
$\Gamma^{(\beta)}=\pi^{-\frac{\beta}{2}N^{2}}\left(\frac{\beta}{2}\right)^{\frac{1}{2}N(N-L)}\prod_{j=1}^{N}\frac{\Gamma(\frac{\beta}{2}j)}{\Gamma(\frac{\beta}{2}(j+L)}$ (7)
@@
$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{\Gamma_{N}}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\\$ (8)
@@
$\Gamma_{N}=\int\cdots\int\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\prod_{j=1}^{N}d^{2}\lambda_{j}=N!\pi^{N}\prod_{k=1}^{N}\Gamma(k+L)$ (9)
@@ 2. Limit theorems for the complex Ginibre ensemble
where each random variables $\gamma_{k}$ are independent and follows a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$ , with $\alpha_{k}=k$ and $\delta_{k}=1$. The Gamma-Rayleigh distribution is derived in the work of E. Akarawak, I. Adeleke and R. Okafor [1].
The corresponding probability density function of the largest modulus $r_{max}^{(N)}(A)$ is derived from Equation (13).
Taking the logarithm of the cumulative distribution function $P(r_{max}^{(N)}(A)\leq r)$ and then the first derivative with respect to $r$ of $\log\left(P(r_{max}^{(N)}(A)\leq r)\right)$, the probability density funvtion of the spectral radius $r_{max}^{(N)}(A)$ is,
@@ 2.1. Limiting left and right tail distributions of the minimum modulus as $N$ goes to infinity
Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble. The left tail distribution of minimum modulus $r_{min}^{(N)}(A)$ (i.e., the smallest values of $r_{min}^{(N)}(A)$) converges to the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$ as $N$ goes to infinity
More precisely, for $0<r\ll 1$,
$\lim_{N\longrightarrow+\infty}P(r_{min}^{(N)}(A)<r)=1-e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (18)
@@
$\lim_{N\rightarrow+\infty}P\left(R_{N}\leq R^{(N,x)}\right)=\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\gamma\left(k+1,N\left(R^{(N,x)}\right)^{2}\right)}{\Gamma(k+1)}=F_{X}(x)=e^{-e^{-x}}$
@@
$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=\lim_{N\rightarrow+\infty}E_{0}^{(2)}\left(\sqrt{N}r\right)$
@@ 3.1. Limit theorems of the outer and inner edges of the ring
Let $A$ denote a $N\times N$ random matrix from the complex Ginibre ensemble. The limiting left tail probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is the probability density function of the Rayleigh distribution with parameter $\frac{1}{\sqrt{2}}$, as $r$ goes to zero and $N$ goes to infinity. More precisely, for $0<r\ll 1$,
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (34)
@@
The limit of partial product $\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$, as $N$ goes to infinity, has been investigated in Section 2 for the complex Ginibre ensemble, and for $0<r\ll 1$, it is expressed as
@@
The limit of the partial sum is derived as follows. Knowing that, for $0<r\ll 1$,
$\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$
@@
Thus, the limiting left tail probability density function of the minimum modulus for matrices from the complex Ginibre ensemble is
@@
The minimum modulus of a $N\times N$ matrix $G$ from the complex induced Ginibre ensemble, with rectangularity index $L=1$, has a left tail probability density function of the Generalised Gamma distribution with scale parameter $\alpha=1$ and other parameters $\kappa=2$ and $\gamma=4$. The corresponding limiting probability density function, for $0<r\ll 1$, as $N$ goes to infinity, is
$\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=2r^{3}e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (35)
@@
Also, the upper incomplete Gamma function $\Gamma(k,x)$, as $x$ approaches zero and $k\in\mathbb{N}\setminus\{0\}$, is $\Gamma(k,x)=\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$.
For $r$ close to zero, $\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$. This implies that the limit of the partial sum, as $r$ approaches zero and $N$ goes to infinity, is
@@
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L\geq 0$. The left tail distribution of minimum of the moduli $r_{min}^{(N)}(G)$ is the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$ in the limit as $N$ goes to infinity and $r$ approaches zero,
@@
This result is derived using the same methodological approach presented in section 2.1.
For $0<r\ll 1$,
@@
Setting the rectangularity index $L$ to zero, the limiting distribution of the minimum modulus $r_{min}^{(N)}(G)$ for the complex induced Ginibre ensemble corresponds to the limiting distribution of the minimum modulus $r_{min}^{(N)}(A)$ for the complex Ginibre ensemble, i.e., for $0<r\ll 1$,
@@
For $0<r\ll 1$, as $N$ goes to infinity, and for a rectangularity index $L=0$,
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)-O\left(r^{5}\right)$